OpenStudy (anonymous):

If you pay $45 for a dress that has been discounted 15%, what is the original price to the nearest cent? Please show how you got your solution.

OpenStudy (anonymous):

15% is the equivalent of 1.5

OpenStudy (anonymous):

45 * 1.5 = original price

OpenStudy (godlovesme):

@cherri8 do u still need help?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

I just gave you help though...

OpenStudy (anonymous):

I mean I literally just set the entire equation up for you

OpenStudy (godlovesme):

like @thedogkeeper said 15% means 15/100 which = 0.15 so if u multiply 45 by 0.15 u will get 6.75 :)

OpenStudy (anonymous):

I appreciate your help, but I thought 15% is the same as 0.15, not 1.5?

OpenStudy (anonymous):

@Godlovesme That is incorrect

OpenStudy (godlovesme):

yes it's 0.15

OpenStudy (godlovesme):

how :/ @thedogkeeper

OpenStudy (anonymous):

look at the question

OpenStudy (anonymous):

it's been discounted already to get to $45 so the original price would obviously have to be higher

OpenStudy (anonymous):

so no it can't be $6.75

OpenStudy (godlovesme):

ohhhh sorry guys so 6.75 is the discount @thedogkeeper

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

So you would add 6.75 to 45.00 to get the original price? Right?

OpenStudy (anonymous):

correct

OpenStudy (godlovesme):

:) 6.75 + 45 = 51.75 is ur original price @cherri8

OpenStudy (anonymous):

Thanks too the both of you!

OpenStudy (godlovesme):

np :)